OpenStudy (anonymous):

integral x/(x^2-2xcos(a)+1) ?

OpenStudy (anonymous):

I assume that a is just any given constant is that correct?

OpenStudy (anonymous):

So you're taking this integral with respect to the x differential, dx dimension.

OpenStudy (anonymous):

a is a constant

OpenStudy (anonymous):

For the denominator consider the following: \[x^2-2x \cos(a)+1=(x- \cos(a))^2 +1 - \cos(a)^2 =(x- \cos a)^2+ \sin^2 a\]

OpenStudy (anonymous):

That should already simplify your integral by a good bit.

OpenStudy (anonymous):

From there I'd recommend to continue with an u-substitution. I think that's all.

OpenStudy (anonymous):

u substitute for (x−cos(a))?

OpenStudy (anonymous):

exactly that's what I would do, then you can express the numerator in terms of that too, will result in two easy integrals. Since a is any constant, du=dx

OpenStudy (anonymous):

Did that help you to get the answer @atsitpfp ?

OpenStudy (anonymous):

yes, thank you, i've got the right answer :)

OpenStudy (anonymous):

great, well done!